Reports 1-1 of 1 Clear search Modify search
AdV-DAQ (Calibration)
estevez, rolland - 10:27 Tuesday 14 May 2019 (45884) Print this report
Calibration Tests

This morning from 6:27 UTC to 7:55 UTC I performed calibration tests after a few changes we did in calibration scripts:

  1. PRWI configuration is now with BS in LowNoise mode (HighPower before). The SNR of the injected lines on BS remains the same as before since there is just a gain of ~4.9 between Sc_BS_MIR_Z_CORR and Sc_BS_MIR_Z_CORR_LN.
  2. The logic of the CALIBRATION_FAILED state was wrong and has been modified.
  • From 6:27 to 6:57, I tried to run PR-WI noise injections with the new PRWI configuration. CALI node went from CALIBRATING_PRWI to CALIBRATED_PRWI in ~10 seconds but didn't perform any injections. I tried it several times. This behavior had already been seen during the last calibration shift.
  • At 6:57 I tried to stop and start CALI process from VPM, then the PRWI injections worked.
  • From 6:57 to 7:55 I tested different ways to go to CALIBRATION_FAILED state (manually from CALI, down from ITF...) during PRWI injections. Every time it went fine. It stopped the noise injections and then restarted the injections from the last failed one.

The changes we did and tried this morning should prevent from strange behaviors of CALI we saw during last week calibration shift where the noise injections were not performed although CALI went through all the steps without complaining. We should also be able to calibrate BS mirror actuator with the PCal as reference.

Search
×

Warning

×